# SegAssess: Panoramic quality mapping for robust and transferable unsupervised segmentation assessment

### v1

@@ III-A Promptable HQ-SAM backbone
where T means transpose operation, $\sqrt{d_{k}}$ refers to the dimensions of keys, Q{Q}, K{K} and V{V} are query (Q{Q}), key (K{K}) and values (V{V}), respectively. Concretely, in the SA layer, Q{Q}, K{K} and V{V} are $O_{f}$} processed by three parallel linear projections, and in the first CA layer, the Q{Q} is linearly projected $O_{f}$} while K and V are projected from $F$. The settings in the second CA layer is opposite to the former one. The outputs of the first TwoWayAttention are then received by the the second one as inputs. This process updates the tokens ${O_{f}}^{8\times{D_{pr}}}$ to ${{O_{f}}^{\prime}}^{8\times{D_{pr}}}$. Afterwards, a final cross-attention layer performs one last interaction between $O_{f}$$^{\prime}$ and visual features $F$, yielding the feature-infused output tokens ${O_{m}}^{8\times{D_{pr}}}$ and updated visual descriptors ${F^{\prime}}^{{D_{pr}}\times(\frac{H}{16}\times\frac{W}{16})}$.
Then the updated tokens ${O_{m}}^{8\times{D_{pr}}}$ are split back into ${O^{\prime}}^{4\times{D_{pr}}}$ and ${{O_{hq}}^{\prime}}^{4\times{D_{pr}}}$. They are subsequently processed by two different MLP blocks for initial and high-quality assessment predictions, respectively. In terms of initial assessment, the visual descriptors ${F^{\prime}}^{{D_{pr}}\times(\frac{H}{16}\times\frac{W}{16})}$ is upsampled with ratio of 4 to ${F^{\prime}}^{{D_{pr}}\times(\frac{H}{4}\times\frac{W}{4})}$ via two transposed convolution layers and element-wise multiplied with the processed standard tokens ${O^{\prime}}^{4\times{D_{pr}}}$. The result is reshaped and upsampled to the full resolution initial assessment mask ${A_{init}}^{H\times W}$. As to high-quality assessment, ${F_{hq}}^{{D_{pr}}\times\frac{H}{4}\times\frac{W}{4}}$ and ${F^{\prime}}^{{D_{pr}}\times(\frac{H}{4}\times\frac{W}{4})}$ are added and merged by two convolution layers to obtain refined features ${{F_{hq}}^{\prime}}^{{D_{pr}}\times(\frac{H}{4}\times\frac{W}{4})}$ fusing high and low level features. After that, ${{F_{hq}}^{\prime}}^{{D_{pr}}\times(\frac{H}{4}\times\frac{W}{4})}$ and ${{O_{hq}}^{\prime}}^{4\times{D_{pr}}}$ are multiplied, reshaped and upsampled to the high-quality assessment mask ${A_{hq}}^{H\times W}$. Finally the ${A_{init}}^{H\times W}$ and ${A_{hq}}^{H\times W}$ are summed up to produce the coarse PQM output from the backbone, marked as ${A_{1}}^{H\times W}$.
@@ III-B Edge guided compaction branch
The edge guided compaction branch leverages the multi-scale image embeddings $F_{1}$}, $F_{2}$}, $F_{3}$} and $F_{4}$} generated by the four stages of the image encoder. Specifically, the lower-level features ${F_{1}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$, ${F_{2}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$, and ${F_{3}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ are independently upsampled by two transposed convolutions to preliminary edge prediction sideouts ${E_{1}}^{H\times W}$, ${E_{2}}^{H\times W}$ and ${E_{3}}^{H\times W}$, respectively. The highest-level image features ${F_{4}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ are first processed by the Aggregated Semantic Filter (ASF) module (see Section III-C) to enhance relevant semantic context for edge detection, and then similarly upsampled to produce the fourth edge sideout ${E_{4}}^{H\times W}$ via transpose convolutions. These four sideouts are then concatenated channel-wise into hierarchical sideouts ${E_{ms}}^{4\times H\times W}$. Simultaneously, three convolutions with $1\times 1$ kernel size are utilized to process ${E_{4}}^{H\times W}$ and produce a weighting map ${W}^{4\times H\times W}$. The ${E_{ms}}^{4\times H\times W}$ and ${W}^{4\times H\times W}$ are multiplied and added up along the channel dimension to generate the final edge prediction ${E}^{H\times W}$. Given ${E}^{H\times W}$ and coarse assessment ${A_{1}}^{H\times W}$, then final SQA results ${A}^{H\times W}$ is calculated following Eq. 2:
@@ V-A1 Effectiveness of different components
Table I: Impacts of different backbone sizes. FZ means freezing image encoder during training. Compares configurations using SAM-B (FZ: frozen encoder, EA: adapter tuning, Base: fully fine-tuned) and fully fine-tuned SAM-L (Large) and SAM-H (Huge). Metrics include trainable parameters (#Param.), inference speed (FPS), and accuracy (mF1/mIoU) on Train-Val, ZS-MS, and ZS-HS scenarios.
@@ V-A2 Influences of different backbone sizes
Table II: Quantitative comparison for the standard Training-Validation protocol. Compares SegAssess (trained with AMS using 4 model types) against AQSNet (trained conventionally) on the 24 model-source SQA datasets (X-Y, where Y $\in$ DeepLabv3+, HRNet, TransUNet, UNetFormer). SegAssess reports metrics for all 4 PQM classes (TP/FP/TN/FN), while AQSNet reports only FP/FN metrics (highlighted in grey for direct comparison). - indicates scores below 1e-4.
@@
V-B Multi-scenario model validation and performance evaluation
@@
Figure 10: Pearson correlation analysis between dataset statistics and model performance. Heatmaps show correlations for AQSNet (left) and SegAssess (right), considering data characteristics (e.g., class percentages, EIB@3, input mask mIoU) and accuracy metrics (per-class F1/IoU). Calculations were based on results from TableII and Table III.
V-B1 Standard training-validation protocol
To establish baseline performance and validate the fundamental effectiveness of SegAssess, we conduct comprehensive experiments following the standard training-validation protocol on 24 benchmark datasets (described in Section. IV-B), i.e. X-DeepLabv3+, X-HRNet, X-TransUNet, X-UNetFormer. Given that SegAssess implements the four-class Panoramic Quality Mapping (PQM) paradigm (TP/FP/TN/FN), simpler binary error localization methods, such as [36], are not directly comparable due to the incompatibility in both label space and evaluation granularity. Among priors, ASQNet [35], which distinguishes FP and FN errors, represents the closest benchmark, although it lacks the comprehensive assessment capabilities of SegAssess. This section systematically evaluates SegAssess through: (1) intra-dataset performance analysis across 24 datasets with F1 and IoU scores, and (2) comparison against ASQNet under identical experimental conditions.
@@ V-B1 Standard training-validation protocol
Qualitative comparisons presented in Fig. 9 visually corroborate the quantitative results. SegAssess consistently produces accurate PQM maps across different datasets and mask sources, whereas AQSNet shows instability (2nd row 3rd column in Fig. V-B) and frequent failures (the last row 3rd column in Fig. V-B). Furthermore, AQSNet’s independent FP and FN predictions often result in significant spatial overlap (Fig. V-B pink areas), indicating difficulty in distinguishing error types, an issue largely resolved by SegAssess’s unified four-class PQM approach. Overall, under standard training and validation conditions, SegAssess demonstrates superior accuracy, stability, and assessment granularity compared to the prior state-of-the-art.
Table III: Quantitative comparison between SegAssess zero-shot inference (Zero-shot) and AQSNet standard training-validation (Train-Val). Comparable metrics are highlighted by grey.
@@
V-B2 Deployment-oriented zero-shot transferring
A critical requirement for practical SQA frameworks is the ability to generalize to segmentation masks generated by unseen models or possessing characteristics different from the training data. We evaluated this crucial zero-shot (ZS) transferability of SegAssess using the reserved test sets: the six X-OCRNet variants representing unseen model-source (ZS-MS) masks, and the BAQS-raw and WAQS-raw datasets representing unseen human-source (ZS-HS) masks. SegAssess, trained using the AMS strategy (which did not include OCRNet or raw human masks), was applied directly to these datasets without any further fine-tuning. We compared its ZS performance against the performance achieved by AQSNet under its standard Train-Val protocol (see Section V-B1), providing a challenging baseline.
@@ V-B2 Deployment-oriented zero-shot transferring
Similarly, when assessing the unseen human-sourced masks in the ZS-HS scenario (BAQS-raw and WAQS-raw), zero-shot SegAssess again exhibited superior accuracy compared to the trained AQSNet baseline. While AQSNet achieved its best results on these specific datasets (potentially due to its original design focus), SegAssess’s inherent four-class PQM approach provided a more comprehensive and accurate assessment even without prior exposure to these human annotation styles. Interestingly, SegAssess achieved higher overall scores on the ZS-HS datasets compared to the ZS-MS datasets. This could be partly attributed to the specific error distributions in the raw datasets (Fig. 6): BAQS-raw has a relatively high proportion of FN pixels, and WAQS-raw has a relatively high proportion of FP pixels, which might slightly inflate the corresponding class metrics and the overall average scores.
@@
In conclusion, these experiments demonstrate SegAssess’s significant capability for zero-shot transfer to both unseen model architectures and unseen human annotation styles, validating its robustness and practical utility for deployment in real-world SQA applications where the source or quality characteristics of the segmentation mask may not be known a priori.

### v2

@@ III-A Promptable HQ-SAM backbone
where T means transpose operation, $\sqrt{d_{k}}$ refers to the dimensions of keys, Q{Q}, K{K} and V{V} are query (Q{Q}), key (K{K}) and values (V{V}), respectively. Concretely, in the SA layer, Q{Q}, K{K} and V{V} are $O_{f}$} processed by three parallel linear projections, and in the first CA layer, the Q{Q} is linearly projected $O_{f}$} while K{K} and V{V} are projected from $F$. The settings in the second CA layer is opposite to the former one. The outputs of the first TwoWayAttention are then received by the the second one as inputs. This process updates the tokens ${O_{f}}^{8\times{D_{pr}}}$ to ${{O_{f}}^{\prime}}^{8\times{D_{pr}}}$. Afterwards, a final cross-attention layer performs one last interaction between $O_{f}$$^{\prime}$ and visual features $F$, yielding the feature-infused output tokens ${O_{m}}^{8\times{D_{pr}}}$ and updated visual descriptors ${F^{\prime}}^{(\frac{H}{16}\times\frac{W}{16})\times{D_{pr}}}$.
Then the updated tokens ${O_{m}}^{8\times{D_{pr}}}$ are split back into ${O^{\prime}}^{4\times{D_{pr}}}$ and ${{O_{hq}}^{\prime}}^{4\times{D_{pr}}}$. They are subsequently processed by two different MLP blocks for initial and high-quality assessment predictions, respectively. In terms of initial assessment, the visual descriptors ${F^{\prime}}^{(\frac{H}{16}\times\frac{W}{16})\times{D_{pr}}}$ is upsampled with ratio of 4 to ${F^{\prime}}^{(\frac{H}{4}\times\frac{W}{4})\times{D_{pr}}}$ via two transposed convolution layers and element-wise multiplied with the processed standard tokens ${O^{\prime}}^{4\times{D_{pr}}}$. The result is reshaped and upsampled to the full resolution initial assessment mask ${A_{init}}^{4\times H\times W}$. As to high-quality assessment, high-quality image representations ${F_{hq}}^{{D_{pr}}\times\frac{H}{4}\times\frac{W}{4}}$ and reshaped visual descriptors ${F^{\prime}}^{{D_{pr}}\times\frac{H}{4}\times\frac{W}{4}}$ are added and merged by two convolution layers, and then flattened to obtain refined features ${{F_{hq}}^{\prime}}^{(\frac{H}{4}\times\frac{W}{4})\times{D_{pr}}}$, which fuses high and low level features. After that, ${{F_{hq}}^{\prime}}^{(\frac{H}{4}\times\frac{W}{4})\times{D_{pr}}}$ and ${{O_{hq}}^{\prime}}^{4\times{D_{pr}}}$ are multiplied, reshaped and upsampled to the high-quality assessment mask ${A_{hq}}^{4\times H\times W}$. Finally the ${A_{init}}^{4\times H\times W}$ and ${A_{hq}}^{4\times H\times W}$ are summed up to produce the coarse PQM output from the backbone, marked as ${A_{1}}^{4\times H\times W}$.
@@ III-B Edge guided compaction branch
The edge guided compaction branch leverages the multi-scale image embeddings $F_{1}$}, $F_{2}$}, $F_{3}$} and $F_{4}$} generated by the four stages of the image encoder. Specifically, the lower-level features ${F_{1}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$, ${F_{2}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$, and ${F_{3}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ are independently upsampled by two transposed convolutions to preliminary edge prediction sideouts ${E_{1}}^{H\times W}$, ${E_{2}}^{H\times W}$ and ${E_{3}}^{H\times W}$, respectively. The highest-level image features ${F_{4}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ are first processed by the Aggregated Semantic Filter (ASF) module (see Section III-C) to enhance relevant semantic context for edge detection, and then similarly upsampled to produce the fourth edge sideout ${E_{4}}^{H\times W}$ via transpose convolutions. These four sideouts are then concatenated channel-wise into hierarchical sideouts ${E_{ms}}^{4\times H\times W}$. Simultaneously, three convolutions with $1\times 1$ kernel size are utilized to process ${E_{4}}^{H\times W}$ and produce a weighting map ${W}^{4\times H\times W}$. The ${E_{ms}}^{4\times H\times W}$ and ${W}^{4\times H\times W}$ are multiplied and added up along the channel dimension to generate the final edge prediction ${E}^{H\times W}$. Given ${E}^{H\times W}$ and coarse assessment ${A_{1}}^{4\times H\times W}$, then final SQA results ${A}^{4\times H\times W}$ is calculated following Eq. 2:
@@ V-A1 Effectiveness of different components
Table I: Impacts of different backbone sizes. FZ means freezing image encoder during training. This table compares configurations using SAM-B (FZ: frozen encoder, EA: adapter tuning, Base: fully fine-tuned) and fully fine-tuned SAM-L (Large) and SAM-H (Huge). Metrics include trainable parameters (#Param.), inference speed (FPS), and accuracy (mF1/mIoU) on Train-Val, ZS-MS, and ZS-HS scenarios.
@@ V-A2 Influences of different backbone sizes
Table II: Quantitative comparison for the standard Training-Validation protocol. This table compares SegAssess (trained with AMS using 4 model types) against AQSNet (trained conventionally) on the 24 model-source SQA datasets (X-Y, where Y $\in$ DeepLabv3+, HRNet, TransUNet, UNetFormer). SegAssess reports metrics for all 4 PQM classes (TP/FP/TN/FN), while AQSNet reports only FP/FN metrics (highlighted in grey for direct comparison). - indicates scores below 1e-4.
@@
V-B Performance Comparison with the State-of-the-Art PEL methods
@@
Figure 10: Pearson correlation analysis between dataset statistics and model performance. Heatmaps show correlations for AQSNet (left) and SegAssess (right), considering data characteristics (e.g., class percentages, EIB@3, input mask mIoU) and accuracy metrics (per-class F1/IoU). Calculations were based on results from TableII and Table IV.
To validate the fundamental effectiveness of our proposed Panoramic Quality Mapping (PQM) paradigm, it is imperative to benchmark it against the state-of-the-art Pixel-level Error Localization (PEL) paradigm. Given that SegAssess implements the four-class Panoramic Quality Mapping (PQM) paradigm (TP/FP/TN/FN), simpler binary error localization methods, such as [21], are not directly comparable due to the incompatibility in both label space and evaluation granularity. Among priors, ASQNet [20], which distinguishes FP and FN errors, represents the closest benchmark, although it lacks the comprehensive assessment capabilities. Therefore, in this section, we conduct extensive experiments to systematically evaluates SegAssess through: (1) intra-dataset performance analysis with F1 and IoU scores following the standard training-validation protocol on 24 benchmark datasets (described in Section. IV-B), i.e. X-DeepLabv3+, X-HRNet, X-TransUNet, X-UNetFormer, and (2) comparison against ASQNet under identical experimental conditions.
@@
Qualitative comparisons presented in Fig. 9 visually corroborate the quantitative results. SegAssess consistently produces accurate PQM maps across different datasets and mask sources, whereas AQSNet shows instability (2nd row 3rd column in Fig. 9) and frequent failures (the last row 3rd column in Fig. 9). Furthermore, AQSNet’s independent FP and FN predictions often result in significant spatial overlap (Fig. 9 pink areas), indicating difficulty in distinguishing error types, an issue largely resolved by SegAssess’s unified four-class PQM approach. Overall, under standard training and validation conditions, SegAssess demonstrates superior accuracy, stability, and assessment granularity compared to the prior state-of-the-art.
V-C Fair Comparison across Different Pixel-wise SQA Granularities
Following the direct comparison with the advanced PEL method AQSNet, we now broaden the analysis to facilitate a fair comparison with PEL methods of varying granularity. Specifically, we introduce a common ’correct/wrong’ binary evaluation framework to directly compare against the paradigm of [21]. By converting the outputs of SegAssess (TP/TN vs. FP/FN) and AQSNet (TP/TN vs. the rest) into this shared binary space, we can conduct a fair and direct side-by-side evaluation across three distinct pixel-wise SQA paradigms.
The quantitative results presented in Table III clearly validate our hypothesis: finer assessment granularity translates to superior performance, even when evaluated under a simplified binary scheme. The method from [21], representing the coarsest ’correct/wrong’ paradigm, serves as a foundational baseline and, as expected, yields the lowest accuracy. AQSNet, by introducing a more nuanced FP/FN distinction, generally achieves significantly better error localization. However, as noted previously, its dual-mask prediction strategy is prone to model collapse on challenging datasets (e.g., GID variants). Ultimately, SegAssess, operating under the most comprehensive four-class PQM paradigm, consistently and stably outperforms both PEL networks across all datasets. This demonstrates that by learning to distinguish all four quality categories, the model develops a more robust and accurate underlying representation of segmentation quality, a superiority that persists even when its output is simplified. These findings verify the discussion in Section V-B.
Table III: Quantitative comparison across different pixel-wise SQA paradigms with the standard Training-Validation protocol. SegAssess and AQSNet results are converted to ’correct/wrong’ binary formant to measure accuracy.
Dataset [21] (Train-Val) AQSNet (Train-Val) SegAssess (Train-Val)
mF1 mIoU mF1 mIoU mF1 mIoU
Inria-DeepLabv3+ 60.31 54.46 66.52 58.03 71.85 61.92
Inria-HRNet 63.77 56.75 69.17 60.18 91.76 84.82
Inria-TransUNet 64.62 68.81 59.92 59.90 71.39 61.73
Inria-UNetFormer 63.95 56.78 69.39 60.09 73.22 63.35
CrowdAI-DeepLabv3+ 70.71 62.21 71.25 62.25 75.41 65.81
CrowdAI-HRNet 80.43 71.03 86.56 78.31 83.40 74.13
CrowdAI-TransUNet 64.22 57.15 71.47 62.16 77.19 67.52
CrowdAI-UNetFormer 82.72 73.59 86.59 78.34 83.48 74.23
DeepGlobe-DeepLabv3+ 64.64 58.00 70.17 61.77 73.15 64.21
DeepGlobe-HRNet 69.82 61.86 68.80 60.69 73.03 64.09
DeepGlobe-TransUNet 69.23 61.43 70.79 63.20 73.10 64.12
DeepGlobe-UNetFormer 71.06 62.78 74.65 65.64 76.36 67.07
GID-DeepLabv3+ 55.97 52.99 49.83 49.60 67.75 60.83
GID-HRNet 61.05 55.90 0.91 0.46 68.71 61.76
GID-TransUNet 55.91 52.62 1.23 0.63 70.72 63.51
GID-UNetFormer 52.38 50.88 49.75 49.50 69.42 62.15
BAQS-DeepLabv3+ 67.69 60.52 70.68 62.65 72.26 64.01
BAQS-HRNet 64.98 58.55 0.94 0.46 70.20 62.23
BAQS-TransUNet 69.72 62.04 71.53 63.37 71.02 62.83
BAQS-UNetFormer 67.80 60.56 49.70 49.41 71.46 65.79
WAQS-DeepLabv3+ 71.58 63.48 72.79 64.39 73.84 64.84
WAQS-HRNet 61.05 55.90 0.95 0.48 69.78 61.70
WAQS-TransUNet 63.66 57.59 1.01 0.51 70.73 62.47
WAQS-UNetFormer 65.79 59.07 49.74 49.49 70.57 62.32
V-D Generalization and Zero-Shot Transferability Analysis
Table IV: Quantitative comparison between SegAssess zero-shot inference (Zero-shot) and AQSNet standard training-validation (Train-Val). Comparable metrics are highlighted by grey.
@@ V-D Generalization and Zero-Shot Transferability Analysis
Table V: Quantitative evaluation of SegAssess’s zero-shot transferability to unseen image/object domains.
Source Domain Target datasets F1TP IoUTP F1FP IoUFP F1TN IoUTN F1FN IoUFN mF1 mIoU
Cross image domain (Inria building $\rightarrow$ CrowdAI building):
Inria CrowdAI-DeepLabv3+ 79.99 66.72 17.77 9.76 93.11 87.12 15.59 8.46 51.62 43.02
CrowdAI-HRNet 80.87 67.96 22.53 12.76 93.05 87.01 33.44 20.14 57.47 46.97
CrowdAI-TransUNet 80.21 67.04 19.90 11.06 93.18 87.24 16.74 9.15 52.51 43.62
CrowdAI-UNetFormer 80.86 67.94 27.91 16.24 93.07 87.05 32.98 19.78 58.70 47.75
CrowdAI-OCRNet 79.87 66.56 17.41 9.55 93.10 87.10 15.61 8.48 51.50 42.92
Cross image and object domain (BAQS building $\rightarrow$ WAQS water):
BAQS WAQS-DeepLabv3+ 77.29 64.65 26.07 15.06 98.61 97.25 4.29 2.20 51.56 44.79
WAQS-HRNet 76.42 63.62 18.85 10.49 98.63 97.29 4.67 2.41 49.64 43.45
WAQS-TransUNet 76.55 63.70 22.15 12.53 98.60 97.24 4.02 2.06 50.33 43.88
WAQS-UNetFormer 76.88 64.06 20.28 11.36 98.60 97.25 4.22 2.18 50.00 43.71
WAQS-OCRNet 77.48 64.79 23.06 13.12 98.61 97.28 3.81 1.96 50.74 44.28
WAQS-raw 77.17 64.20 59.71 42.78 98.00 96.09 7.05 3.72 60.48 51.70
@@
A critical measure of a practical SQA framework is its ability to generalize to conditions unseen during training. In this section, we rigorously evaluate the transferability of SegAssess from two critical and progressively challenging perspectives. First, we assess its ability to handle segmentation masks from new sources (both model- and human-generated) while remaining within the learned image and object domains. Second, we probe the boundaries of its capabilities by testing its more demanding zero-shot performance when transferring to entirely new image domains and object categories. This two-tiered analysis provides a comprehensive picture of the model’s robustness and practical deployment potential.
V-D1 Zero-shot transferring to unseen masks
A critical requirement for practical SQA frameworks is the ability to generalize to segmentation masks generated by unseen models or possessing characteristics different from the training data. We evaluated this crucial zero-shot (ZS) transferability of SegAssess using the reserved test sets: the six X-OCRNet variants representing unseen model-source (ZS-MS) masks, and the BAQS-raw and WAQS-raw datasets representing unseen human-source (ZS-HS) masks. SegAssess, trained using the AMS strategy (which did not include OCRNet or raw human masks), was applied directly to these datasets without any further fine-tuning. We compared its ZS performance against the performance achieved by AQSNet under its standard Train-Val protocol (see Section V-B), providing a challenging baseline.
@@ V-D1 Zero-shot transferring to unseen masks
Similarly, when assessing the unseen human-sourced masks in the ZS-HS scenario (BAQS-raw and WAQS-raw), zero-shot SegAssess again exhibited superior accuracy compared to the trained AQSNet baseline. While AQSNet achieved its best results on these specific datasets (potentially due to its original design focus), SegAssess’s inherent four-class PQM approach provided a more comprehensive and accurate assessment even without prior exposure to these human annotation styles. Interestingly, SegAssess achieved higher overall scores on the ZS-HS datasets compared to the ZS-MS datasets. This could be partly attributed to the specific error distributions in the raw datasets (Fig. 6): BAQS-raw and WAQS-rawhas have a relatively high proportion of FN and FP pixels, respectively, which might slightly inflate the corresponding class metrics and the overall average scores.
@@
In conclusion, these experiments demonstrate SegAssess’s significant capability for zero-shot transfer to both unseen model-or human-source masks, validating its robustness and practical utility for deployment in real-world SQA applications where the source or quality characteristics of the segmentation mask may not be known a priori.
V-D2 Zero-shot transferring to unseen image/object domains
Building upon the analysis of mask-source transferability, we now investigate a more stringent test of generalization: the model’s performance when transferred to entirely new image or object domains. This directly probes the boundaries of the model’s learned representations. As detailed in Table V we conducted two cross-domain experiments: one transferring from the CrowdAI (building) domain to the Inria (building) domain to assess cross-image-domain performance, and another transferring from the BAQS (building) domain to the WAQS (water) domain to evaluate cross-image-and-object-domain performance.
[…]
In summary, this two-tiered analysis clarifies the generalization capabilities and current limitations of SegAssess. It demonstrates remarkable robustness in zero-shot transfer to unseen mask sources within a consistent image and object domain (as shown in Section V-D1). However, transferring across domains remains a significant challenge, particularly for the identification of minority error classes (FP/FN). This distinction defines the current operational boundaries of SegAssess and highlights critical directions for future work, as discussed in Section VI.